OpenStudy (anonymous):

@tkhunny The social convenor has 12 volunteers to work at a school dance. Each dance requires 2 volunteers at the door, 4 volunteers on the floor, and 6 floaters. Joe and Jim have not volunteered before, so the social convenor does not want to assign them to work together. In how many ways can the volunteers be assigned? please check my answer

OpenStudy (anonymous):

Indirect method without any restriction: 12C2*10C4*6C6=13860 Combinations with both of them working together at the Door: 1*10C4*6C6=210 At the Floor: 10C2*8C2*6C6=1260 As Floaters: 10c2*8C2*4C4=3150 =210+1260+3150=4620 No restriction - What you Dont Want= 13860-4620=9240

OpenStudy (anonymous):

Does my answer make sense, or have i missed something?

OpenStudy (anonymous):

@tkhunny any idea yet?

OpenStudy (tkhunny):

I tend to be horrible at counting, but the ONLY thing I find wrong with your presentation is the 8C2 in the "Floater" section. s/b 8C4 It's just a typo, since you have the right total.

OpenStudy (anonymous):

oh my bad yes its a typo, thanks :)